Find (a) such that 2x^2a[8x^(5a-4)+5x^(2a+1)]=16x^31+10x^21 Please explain how you got your answer
Answer by Boreal(15235)

2x^2a[8x^(5a-4)+5x^(2a+1)]=16x^31+10x^21
16x^7a-4 + 10x^4a+1
Add exponents when distributing
Then set the exponents equal to the appropriate coefficient of the variable, 16x^ and 10x^
7a-4=31
7a=35
a=5
4a+1=21
4a=20
a=5